An architect designes a rectangular flower garden such that the width is exactly two-thirds of the length. IF 260 feet of antique picket fencing are to be used to enclose the garden, find the dimensions of the garden.
Answer by mananth(16946)

An architect designes a rectangular flower garden such that the width is exactly two-thirds of the length. IF 260 feet of antique picket fencing are to be used to enclose the garden, find the dimensions of the garden.
Let length be x
Width = 2/3 x
2x+4/3 x=260
6x+4x/3 =260 LCD
10x/3 =260
10x=780
X=78 feet
Width = 52 feet